# A General Control-Theoretic Approach for Reinforcement Learning: Theory and Algorithms

### v1

@@ Abstract
We devise a control-theoretic reinforcement learning approach to support direct learning of the optimal policy. We establish theoretical properties of our approach and derive an algorithm based on a specific instance of this approach. Our empirical results demonstrate the significant benefits of our approach.
Keywords: Reinforcement Learning, Control Theory, Policy Gradient Methods.
1 Introduction
@@
For many years now, reinforcement learning (RL) has succeeded in solving a wide variety of decision-making problems and control for robotics [1, 2, 3, 4, 5]. Generally speaking, model-free methods [6, 7] often suffer from high sample complexity that can require an inordinate amount of samples, making them unsuitable for robotic applications where collecting large amounts of data is time-consuming, costly and potentially dangerous for the system and its surroundings [8, 9, 10, 11, 12]. On the other hand, model-based RL methods have been successful in demonstrating significantly reduced sample complexity and in outperforming model-free approaches for various decision making under uncertainty problems (see, e.g., [13, 14]). However, such model-based approaches can suffer from the difficulty of learning an appropriate model and from worse asymptotic performance than model-free approaches due to model bias from inherently assuming the learned system dynamics model accurately represents the true system environment (see, e.g., [15, 16, 17]).
@@
In this paper we propose a novel form of RL that seeks to directly learn an optimal control policy for a general underlying (unknown) dynamical system and to directly apply the corresponding learned optimal control policy within the dynamical system. This general approach is in strong contrast to many traditional model-based RL methods that, after learning the system dynamics model which is often of high complexity and dimensionality, then use this system dynamics model to compute an approximate solution of a corresponding (stochastic) dynamic programming problem, often applying model predictive control (see, e.g., [18]). Our control-based RL (CBRL) approach instead directly learns the unknown parameters that derive, through control-theoretic means, an optimal control policy function from a family of control policy functions, often of much lower complexity and dimensionality, from which the optimal control policy is directly obtained.
@@
The theoretical foundation and analysis of our CRBL approach is presented within the context of a general Markov decision process (MDP) framework that extends the family of policies associated with the classical Bellman operator to a family of control-policy functions mapping a vector of (unknown) parameters from a corresponding parameter set to a control policy which is optimal under those parameters, and that extends the domain of these control policies from a single state to span across all (or a large subset of) states, with the (unknown) parameter vector encoding global and local information that needs to be learned. Within the context of this MDP framework and our general CBRL approach, we establish theoretical results on convergence and optimality with respect to (w.r.t.) a CBRL contraction operator, analogous to the Bellman operator.
@@
The basic idea of learning a form of a parameterized policy within an MDP framework to reduce sample complexity is not a new idea. One popular approach along these lines concerns policy gradient methods [19, 20], where gradient ascent of the value function in a space of policies is used, possibly together with projection to obtain an optimal policy [21]. These ideas have been further refined in neural network based policy optimization approaches (see, e.g., [21]) such as TRPO [22] and PPO [23]. It is important to note, however, that our CBRL approach is fundamentally different in several important ways. More specifically, we exploit control-theoretic methods to derive the optimal policy in terms of estimates of a few (global) parameters and directly learn these unknown parameters in an iterative manner based on observations from applying the optimal control policy of the current estimate of parameters. Meanwhile, within the context of our general CBRL approach, we establish a new control-policy-parameter gradient ascent theorem, analogous to the standard policy gradient theorem, together with a corresponding gradient ascent algorithm based on the (unknown) parameter vector.
@@
With its foundation being optimal control, our CBRL approach is particularly suitable for dynamical systems and therefore the control of robotic systems. In addition to its established theoretical properties, numerical experiments of various simple robotics tasks demonstrate the effectiveness of our CBRL approach in reducing the number of samples needed, which is a key requirement for the application of learning and decision-making algorithms in real robotic systems.
@@
We first present our general CBRL approach and then adapt this approach to a particular control-theoretic framework. Empirical results are presented next, followed by a discussion of limitations and future work. The appendix provides all proofs, experimental details, and additional results.
2 Control-Based Reinforcement Learning Approach
@@
Consider a standard RL framework (see, e.g., [20, 24]) in which a decision-making agent interacts with a stochastic environment modeled as an MDP defined over a set of states ${{\mathbb{X}}}$, a set of actions ${\mathbb{U}}$, a transition probability kernel $\mathbb{P}$, and a reward function ${r}$ mapping state-action pairs to a bounded subset of ${\mathbb{R}}$. We denote a probability distribution on ${\mathbb{U}}$ as ${\mathbb{W}}$ (i.e., ${\mathbb{W}}$ is an element of $[0,1]^{{\mathbb{U}}}$ whose values sum to $1$). The agent seeks to determine a policy $\pi(\cdot|x):{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ comprising the sequence of control actions $u_{1},u_{2},\ldots$ that maximizes the corresponding discounted infinite-horizon stochastic dynamic programming (sDP) formulation expressed as
@@
where $x_{t}\in{{\mathbb{X}}}$ denotes the state of the system at time $t\in{\mathbb{Z}}_{+}$, $u_{t}\in{\mathbb{U}}$ the control action at time $t$, $\gamma\in(0,1)$ the discount factor, and $\mathbb{E}_{\mathbb{P}}:=\mathbb{E}_{x_{t+1}\sim\mathbb{P}(\cdot|x_{t},u)}$ defines expectation w.r.t. $\mathbb{P}(\cdot|x_{t},u)$. The stationary policy $\pi(\cdot|x):{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ defines a distribution of available control actions given the current state $x$, which reduces to a deterministic policy $\pi(\cdot|x):{{\mathbb{X}}}\rightarrow{\mathbb{U}}$ when the conditional distribution renders a constant action for state $x$; with abuse of notation, we write policy $\pi(x)$. Let ${{\mathbb{Q}}}({{\mathbb{X}}}\times{\mathbb{U}})$ denote the space of bounded real-valued functions over ${{\mathbb{X}}}\times{\mathbb{U}}$ with supremum norm, let $\Pi$ denote the set of all stationary policies, and let $\pi^{\ast}(x)$ denote the optimal stationary policy.
@@
Our CBRL approach consists of exploiting control-theoretic methods to solve the general sDP formulation (1) in terms of the unknown system and control variables, and directly learning these unknown parameters over time through a corresponding iterative solution process. In particular, our general approach considers two key ideas: (i) extending the family of policies $\Pi$ associated with classical RL to a family of control-policy functions that map a parameter vector $p$ from a parameter set to a control policy that is optimal under $p$; and (ii) extending the domain of these control policies from a single state to span across all states in ${{\mathbb{X}}}$, with the parameter vector $p$ encoding global and local information (e.g., gravity and mass) that needs to be learned. More formally, let ${{\mathbb{S}}}$ denote the set of parameter vectors, and ${\mathbb{F}}$ the family of control-policy functions whose elements comprise functions ${\mathcal{G}}:p\mapsto{\mathcal{F}}_{p}$ that map a parameter vector $p\in{{\mathbb{S}}}$ to a control policy ${\mathcal{F}}_{p}:{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ that is optimal under $p$. For any $p\in{{\mathbb{S}}}$, the control-policy function ${\mathcal{G}}(\cdot)\in{\mathbb{F}}$ identifies a specific control policy ${\mathcal{F}}_{p}\in\Pi$ that provides the best rewards in expectation across all states $x\in{{\mathbb{X}}}$ from among all control-policy functions in ${\mathbb{F}}$. Such control policy mappings ${\mathcal{G}}:{{\mathbb{S}}}\rightarrow\Pi$ derive optimal control policies from vectors in the parameter set ${{\mathbb{S}}}$ through control-theoretic methods, with the goal of learning the unknown parameter vector $p\in{{\mathbb{S}}}$ through a corresponding iterative solution process. Hence, once the parameter vector is learned, the desired optimal control policy is realized.
@@
In the remainder of this section, we first establish theoretical properties of our CBRL approach and then consider an algorithm to compute solutions based on our CBRL approach and its properties.
@@
Convergence and Optimality. We consider an analog of the Bellman operator and related theoretical results on convergence and optimality. For each $Q$-function $q(x,u)\in{{\mathbb{Q}}}({{\mathbb{X}}}\times{\mathbb{U}})$, define a generic CBRL function $\tilde{q}:{{\mathbb{X}}}\times\Pi\rightarrow{\mathbb{R}}$ as $\tilde{q}(x,{\mathcal{F}}_{p}):=q(x,{\mathcal{F}}_{p}(x))$ where the control policy ${\mathcal{F}}_{p}$ is obtained from the control-policy function ${\mathcal{G}}$ derived from the parameter vector $p\in{{\mathbb{S}}}$, and thus evidently $\tilde{q}(x,{\mathcal{F}}_{p})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ with ~$\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ denoting the space of bounded real-valued functions over ${{\mathbb{X}}}\times\Pi$ with supremum norm. We then define our CBRL operator ${{\mathbf{T}}}$ on ~$\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ as
$({{\mathbf{T}}}\tilde{q})(x,{\mathcal{F}}_{p}):=\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\Big{[}r(x,{\mathcal{F}}_{p}(x))+\gamma\sup_{p^{\prime}\in{{\mathbb{S}}}}\tilde{q}(y,{\mathcal{F}}_{p^{\prime}})\Big{]}.$ (2)
@@
Our approach comprises a progression from the family ${\mathbb{F}}$ to the control-policy functions ${\mathcal{G}}\in{\mathbb{F}}$ to the control policies ${\mathcal{F}}_{p}\in\Pi$ to the control action ${\mathcal{F}}_{p}(x)\in{\mathbb{W}}$, upon applying ${\mathcal{F}}_{p}$ in state $x\in{{\mathbb{X}}}$. We next introduce an important assumption on the richness of the policy family ${\mathbb{F}}$ within this progression.
Assumption 2.1.
@@
There exist a policy function ${\mathcal{G}}$ in the family ${\mathbb{F}}$ and a unique vector $p^{\prime}$ in the parameter set ${{\mathbb{S}}}$ such that, for any state $x\in{{\mathbb{X}}}$, $\pi^{\ast}(x)={\mathcal{F}}_{p^{\prime}}(x)={\mathcal{G}}(p^{\prime})(x)$.
@@
Assumption 2.1 says that ${\mathbb{F}}$ is rich enough to include a global policy that coincides with the Bellman operator for each state. We then have the following formal result on the convergence of the operator ${{\mathbf{T}}}$ of our CBRL approach and the optimality of this convergence w.r.t. the Bellman equation [20, 24].
Lemma 2.1.
@@
For any $\gamma\in(0,1)$, operator ${{\mathbf{T}}}$ in (2) is a contraction in the supremum norm. Suppose Assumption 2.1 holds for the family of policy functions ${\mathbb{F}}$ and its parameter set ${{\mathbb{S}}}$. Then, our CBRL approach with contraction operator ${{\mathbf{T}}}$ achieves the same optimal outcome as the Bellman operator.
@@
From the first part of Lemma 2.1 and for any function $\tilde{q}\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$, the iterations ${{\mathbf{T}}}^{t}(\tilde{q})$ converge as $t\rightarrow\infty$ to $\tilde{q}^{\prime}(x,{\mathcal{F}}_{p})$, the unique fixed point of the contraction operator ${{\mathbf{T}}}$, and $\tilde{q}^{\prime}(x,{\mathcal{F}}_{p})$ satisfies
$\tilde{q}^{\prime}(x,{\mathcal{F}}_{p})=\sum_{y\in{{\mathbb{X}}}}\mathbb{P}_{{\mathcal{F}}_{p}(x)}(x,y)\Big{[}r(x,{\mathcal{F}}_{p}(x))+\gamma\sup_{p^{\prime}\in{{\mathbb{S}}}}\tilde{q}^{\prime}(y,{\mathcal{F}}_{p^{\prime}})\Big{]}.$ (3)
@@
The second part of Lemma 2.1 implies that, under the contraction operator ${{\mathbf{T}}}$ and Assumption 2.1, our CBRL approach is optimal by realizing the same unique fixed point of the Bellman operator. We note, however, that this optimality is achieved with great reduction in sample complexity due in part to another important difference with standard RL, namely the search of our CBRL approach is across all states $x\in{{\mathbb{X}}}$ to find a single optimal parameter vector $p$ that derives a single optimal control-policy function ${\mathcal{G}}\in{\mathbb{F}}$ which coincides with the Bellman equation for each state. Iterations w.r.t. the operator ${{\mathbf{T}}}$ of our CBRL approach then consist of improving the estimates of the parameter vector $p$ while applying the optimal control-policy function ${\mathcal{G}}$ derived from the current estimate of $p$.
@@
Before addressing this iterative process within the context of policy gradient methods, we first relax the above conditions to consider families that do not satisfy Assumption 2.1 and policy functions that map a parameter vector to a control policy whose domain spans across a large subset of states (as opposed to all states) in ${{\mathbb{X}}}$ and renders asymptotically optimal (as opposed to globally optimal) rewards. Supposing ${\mathbb{F}}$ satisfies Assumption 2.1, consider a sequence of less rich families ${\mathbb{F}}_{1}\subset\cdots\subset{\mathbb{F}}_{k}\subset{\mathbb{F}}$ of policy functions ${\mathcal{G}}^{(i)}\in{\mathbb{F}}_{i}$ derived from parameter vectors of the corresponding sets ${{\mathbb{S}}}_{i}$, and define the operators ${{\mathbf{T}}}_{i}:\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})\rightarrow\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})$ as in (2) for any function $\tilde{q}_{i}(x,{\mathcal{F}}_{p}^{(i)})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})$, $i\in[k]$. Then, from Lemma 2.1, each operator ${{\mathbf{T}}}_{i}$ under parameter set ${{\mathbb{S}}}_{i}$ is a contraction in the supremum norm and converges to the unique fixed point $\tilde{q}_{i}^{\prime}(x,{\mathcal{F}}_{p}^{(i)})$ that satisfies the corresponding version of (3), for all $x\in{{\mathbb{X}}}$ and $p\in{{\mathbb{S}}}_{i}$, $i\in[k]:=\{1,\ldots,k\}$. We then have the following formal result on the asymptotic optimality of our CBRL approach in such approximate settings.
Theorem 2.1.
@@
One instance of a sequence of families ${\mathbb{F}}_{1}\subset\cdots\subset{\mathbb{F}}_{k-1}\subset{\mathbb{F}}_{k}$ in Theorem 2.1 consists of piecewise-linear control policies of increasing richness (e.g., the class of canonical piecewise-linear functions [25]) w.r.t. finer and finer granularity of the control policy function space converging towards ${\mathbb{F}}$.
@@
Control-Policy-Parameter Gradient Ascent. We now return to our CBRL iterative process in the context of policy gradient methods, building on our foregoing results to establish theoretical results analogous to the standard policy gradient theorem that establishes, for policies parameterized by $\theta$, a connection between the gradient of the value function w.r.t. $\theta$ and the gradient of the policy action w.r.t. $\theta$. This theoretical result does not apply to our CBRL approach because there is no direct connection between the parameter vector $p\in{{\mathbb{S}}}$ and the control policy action. Rather, under CBRL and for any vector $p\in{{\mathbb{S}}}$, the control-policy function ${\mathcal{G}}(p)\in{\mathbb{F}}$ identifies a particular control policy ${\mathcal{F}}_{p}\in\Pi$, and then the control policy applied in any state $x\in{{\mathbb{X}}}$ yields the control action taken.

### v2

@@ Abstract
We devise a control-theoretic reinforcement learning approach to support direct learning of the optimal policy. We establish various theoretical properties of our approach, such as convergence and optimality of our control-theoretic operator, a new control-policy-parameter gradient ascent theorem, and a specific gradient ascent algorithm based on this theorem. As a representative example, we adapt our approach to a particular control-theoretic framework and empirically evaluate its performance on several classical reinforcement learning tasks, demonstrating significant improvements in solution quality, sample complexity, and running time of our control-theoretic approach over state-of-the-art baseline methods.
Introduction
@@
For many years now, reinforcement learning (RL) has been successful in solving a wide variety of decision making under uncertainty problems (Feng et al. 2024; Shen 2024; McMahan et al. 2024; Zheng et al. 2023). Many different RL approaches, with varying degrees of success, have been developed to address such problems (Kaelbling, Littman, and Moore 1996; Szepesvari 2010; Sutton and Barto 2020). Generally speaking, model-free RL methods (Randlov and Alstrom 1998; Mnih et al. 2013) often suffer from high sample complexity that can require an inordinate amount of samples for some problems, making them unsuitable for various applications where collecting large amounts of data is time-consuming, costly and potentially dangerous for the system and its surroundings (Zhang and Tan 2024; Chen, Subramanian, and Paternain 2024; Peng et al. 2023; Dong et al. 2023; Kumar et al. 2020). On the other hand, model-based RL methods have been successful in demonstrating significantly reduced sample complexity and in outperforming model-free approaches for various decision making under uncertainty problems; see, e.g., (Deisenroth and Rasmussen 2011; Meger et al. 2015). However, such model-based approaches can suffer from the difficulty of learning an appropriate model and from worse asymptotic performance than model-free approaches due to model bias from inherently assuming the learned system dynamics model accurately represents the true system environment; see, e.g., (Atkeson and Santamaria 1997; Schneider 1997; Schaal 1997).
@@
In this paper we propose a novel form of RL that seeks to directly learn an optimal control policy for a general underlying (unknown) dynamical system and to directly apply the corresponding learned optimal control policy within the dynamical system. This general approach is in strong contrast to many traditional model-based RL methods that, after learning the system dynamics model which is often of high complexity and dimensionality, then use this system dynamics model to compute an approximate solution of a corresponding (stochastic) dynamic programming problem, often applying model predictive control; see, e.g., (Nagabandi et al. 2018). Our control-based RL (CBRL) approach instead directly learns the unknown parameters that derive, through control-theoretic means, an optimal control policy function from a family of control policy functions, often of much lower complexity and dimensionality, from which the optimal control policy is directly obtained.
@@
The theoretical foundation and analysis of our CRBL approach is presented within the context of a general Markov decision process (MDP) framework that extends the family of policies associated with the classical Bellman operator to a family of control-policy functions mapping a vector of (unknown) parameters from a corresponding parameter set to a control policy which is optimal under those parameters, and that extends the domain of these control policies from a single state to span across all (or a large subset of) states, with the (unknown) parameter vector encoding global and local information that needs to be learned. Within the context of this MDP framework and our general CBRL approach, we establish theoretical results on convergence and optimality with respect to (w.r.t.) a CBRL contraction operator, analogous to corresponding results for the Bellman operator.
@@
The basic idea of learning a form of a parameterized policy within an MDP framework to reduce sample complexity has been studied in the past. One popular approach along these lines concerns policy gradient methods (Sutton et al. 1999; Sutton and Barto 2020), where gradient ascent of the value function in a space of policies is used, possibly together with projection to obtain an optimal policy (Agarwal et al. 2021). These ideas are further refined in neural network based policy optimization approaches (see, e.g., (Agarwal et al. 2021)) such as TRPO (Schulman et al. 2015) and PPO (Schulman et al. 2017). It is important to note, however, that our CBRL approach is fundamentally different in several important ways. More specifically, we exploit control-theoretic methods to derive the optimal policy in terms of estimates of a few (global) parameters and directly learn these unknown parameters in an iterative manner based on observations from applying the optimal control policy for the current estimate of parameters. Meanwhile, within the context of our CBRL approach, we establish a new control-policy-parameter gradient ascent theorem, analogous to the standard policy gradient theorem, together with a corresponding gradient ascent method that comprises an iterative process for directly learning the (unknown) parameter vector.
@@
With its foundation being optimal control, our CBRL approach is particularly suitable for dynamical systems in general and therefore the optimal control of a wide variety of systems. In addition to its established theoretical properties, numerical experiments of various classical decision making under uncertainty tasks demonstrate the effectiveness of our CBRL approach in reducing the number of samples needed, which is a key requirement for the application of learning and decision-making algorithms in real-world systems.
@@
We first present our general CBRL approach and then adapt this approach to a particular control-theoretic framework. Empirical results are presented next, followed by concluding remarks. The appendix provides all proofs, and additional theoretical and experimental details and results.
CBRL Approach
@@
Consider a standard RL framework (see, e.g., (Sutton and Barto 2020; Bertsekas and Tsitsiklis 1996)) in which a decision-making agent interacts with a stochastic environment modeled as an MDP defined over a set of states ${{\mathbb{X}}}$, a set of actions ${\mathbb{U}}$, a transition probability kernel $\mathbb{P}$, and a reward function ${r}$ mapping state-action pairs to a bounded subset of ${\mathbb{R}}$. We denote a probability distribution on ${\mathbb{U}}$ as ${\mathbb{W}}$. The agent seeks to determine a policy $\pi(\cdot|x):{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ comprising the sequence of control actions $u_{1},u_{2},\ldots$ that maximizes the corresponding discounted infinite-horizon stochastic dynamic programming (sDP) formulation expressed as
@@
where $x_{t}\in{{\mathbb{X}}}$ denotes the state of the system at time $t\in{\mathbb{Z}}_{+}$, $u_{t}\in{\mathbb{U}}$ the control action at time $t$, $\gamma\in(0,1)$ the discount factor, and $\mathbb{E}_{\mathbb{P}}:=\mathbb{E}_{x_{t+1}\sim\mathbb{P}(\cdot|x_{t},u_{t})}$ defines expectation w.r.t. the conditional transition probability for the next state $\mathbb{P}(\cdot|x_{t},u_{t})$. The stationary policy $\pi(\cdot|x):{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ defines a distribution of available control actions given the current state $x$, which reduces to a deterministic policy $\pi(\cdot|x):{{\mathbb{X}}}\rightarrow{\mathbb{U}}$ when the conditional distribution renders a constant action for state $x$; with slight abuse of notation, we write policy $\pi(x)$. Let ${{\mathbb{Q}}}({{\mathbb{X}}}\times{\mathbb{U}})$ denote the space of bounded real-valued functions over ${{\mathbb{X}}}\times{\mathbb{U}}$ with supremum norm, and let $\Pi$ denote the set of all stationary policies. From standard RL theory, define the Bellman operator ${{\mathcal{T}}_{B}}$ on ${{\mathbb{Q}}}({{\mathbb{X}}}\times{\mathbb{U}})$ as
${{\mathcal{T}}_{B}}Q(x,u):={r}(x,u)+\gamma\mathbb{E}_{\mathbb{P}}\max_{u^{\prime}\in{\mathbb{U}}}Q(x^{\prime},u^{\prime}),$ (2)
where $x^{\prime}$ denotes the next state upon transition from $x$. Let $Q^{\ast}(x,u)$, $V^{\ast}(x):=\max_{u\in{\mathbb{U}}}Q^{\ast}(x,u)$, and $\pi^{\ast}(x):=\operatorname*{\arg\max}_{u\in{\mathbb{U}}}Q^{\ast}(x,u)$ denote the optimal action-value function, the optimal value function, and the optimal stationary policy, respectively.
@@
Our CBRL approach consists of exploiting control-theoretic methods to solve the general sDP formulation (1) in terms of the unknown system and control variables, and directly learning these unknown parameters over time through a corresponding iterative solution process. In particular, our general approach considers two key ideas: (i) extending the family of policies $\Pi$ associated with classical RL to a family of control-policy functions that map a parameter vector $p$ from a parameter set to a control policy that is optimal under the parameter vector $p$; and (ii) extending the domain of these control policies from a single state to span across all states in ${{\mathbb{X}}}$, with the parameter vector $p$ encoding both global and local information (e.g., gravity and mass) that is unknown and needs to be learned.
More formally, let ${{\mathbb{S}}}$ denote a subset of a metric space serving as the set of parameter vectors, and ${\mathbb{F}}$ the family of control-policy functions whose elements comprise surjective functions ${\mathcal{G}}:p\mapsto{\mathcal{F}}_{p}$ that map a parameter vector $p\in{{\mathbb{S}}}$ to a control policy ${\mathcal{F}}_{p}:{{\mathbb{X}}}\rightarrow{\mathbb{W}}$ that is optimal under $p$. For any $p\in{{\mathbb{S}}}$, the control-policy function ${\mathcal{G}}(\cdot)\in{\mathbb{F}}$ identifies a specific control policy ${\mathcal{F}}_{p}\in\Pi$ that provides the best expected cumulative discounted reward across all states $x\in{{\mathbb{X}}}$ from among all control-policy functions in ${\mathbb{F}}$. Such control policy mappings ${\mathcal{G}}:{{\mathbb{S}}}\rightarrow\Pi$ derive optimal control policies from vectors in the parameter set ${{\mathbb{S}}}$ through control-theoretic methods, with the goal of learning the unknown parameter vector $p\in{{\mathbb{S}}}$ through a corresponding iterative solution process. Hence, once the parameter vector is learned with sufficient accuracy, the desired optimal control policy is realized.
@@
In the remainder of this section, we first establish theoretical properties of our CBRL approach and then consider a new form of control-policy-parameter gradient ascent methods and corresponding theoretical properties.
@@
Convergence and Optimality. We consider an analog of the Bellman operator ${{\mathcal{T}}_{B}}$ and related theoretical results on convergence and optimality. For each $Q$-function $q(x,u)\in{{\mathbb{Q}}}({{\mathbb{X}}}\times{\mathbb{U}})$, define the generic CBRL function $\tilde{q}:{{\mathbb{X}}}\times\Pi\rightarrow{\mathbb{R}}$ as $\tilde{q}(x,{\mathcal{F}}_{p}):=q(x,{\mathcal{F}}_{p}(x))$ where the control policy ${\mathcal{F}}_{p}$ is obtained from the control-policy function ${\mathcal{G}}$ derived from the parameter vector $p\in{{\mathbb{S}}}$, and thus $\tilde{q}(x,{\mathcal{F}}_{p})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ with ~$\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ denoting the space of bounded real-valued functions over ${{\mathbb{X}}}\times\Pi$ with supremum norm. We then define our CBRL operator ${{\mathbf{T}}}$ on ~$\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$ as
$({{\mathbf{T}}}\tilde{q})(x,{\mathcal{F}}_{p}):=\sum_{y\in{{\mathbb{X}}}}$ $\;\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p})\;\;\times$
$\Big{[}r(x,{\mathcal{F}}_{p}(x))+\gamma\sup_{p^{\prime}\in{{\mathbb{S}}}}\tilde{q}(y,{\mathcal{F}}_{p^{\prime}})\Big{]}.$ (3)
@@
Our approach comprises a sequence of steps from the family ${\mathbb{F}}$ to the control-policy functions ${\mathcal{G}}\in{\mathbb{F}}$ to the control policies ${\mathcal{F}}_{p}\in\Pi$ to the control action ${\mathcal{F}}_{p}(x)\in{\mathbb{W}}$, upon applying ${\mathcal{F}}_{p}$ in state $x\in{{\mathbb{X}}}$. We next introduce an important assumption on the richness of the family ${\mathbb{F}}$ within this sequence of steps.
Assumption 1.
@@
There exist a policy function ${\mathcal{G}}$ in the family ${\mathbb{F}}$ and a unique vector $p^{\ast}$ in the parameter set ${{\mathbb{S}}}$ such that, for any state $x\in{{\mathbb{X}}}$, $\pi^{\ast}(x)={\mathcal{F}}_{p^{\ast}}(x)={\mathcal{G}}(p^{\ast})(x)$.
@@
Assumption 1 says that ${\mathbb{F}}$ is rich enough to include a global policy that coincides with the Bellman operator for each state. We then have the following formal result on the convergence of the operator ${{\mathbf{T}}}$ of our CBRL approach and the optimality of this convergence w.r.t. the Bellman equation (Sutton and Barto 2020; Bertsekas and Tsitsiklis 1996).
Theorem 1.
@@
(a) For any $\gamma\in(0,1)$, the operator ${{\mathbf{T}}}$ in (CBRL Approach) is a contraction in the supremum norm. (b) Suppose Assumption 1 holds for the family of policy functions ${\mathbb{F}}$ and its parameter set ${{\mathbb{S}}}$. Then, our CBRL approach with contraction operator ${{\mathbf{T}}}$ achieves the same asymptotically optimal outcome as that of the Bellman operator ${{\mathcal{T}}_{B}}$.
@@
From Theorem 1(a), for any function $\tilde{q}\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi)$, the iterations ${{\mathbf{T}}}^{t}(\tilde{q})$ converge as $t\rightarrow\infty$ to $\tilde{q}^{\prime}(x,{\mathcal{F}}_{p})$, the unique fixed point of the operator ${{\mathbf{T}}}$, and $\tilde{q}^{\prime}(x,{\mathcal{F}}_{p})$ satisfies
$\tilde{q}^{\prime}(x,{\mathcal{F}}_{p})\!=\!\!\sum_{y\in{{\mathbb{X}}}}\mathbb{P}(y\,|\,x,{\mathcal{F}}_{p})\Big{[}r(x,{\mathcal{F}}_{p}(x))+\gamma\sup_{p^{\prime}\in{{\mathbb{S}}}}\tilde{q}^{\prime}(y,{\mathcal{F}}_{p^{\prime}})\Big{]}.$ (4)
@@
Theorem 1(b) implies that, under the contraction operator ${{\mathbf{T}}}$ and Assumption 1, our CBRL approach is optimal by realizing the same unique fixed point of the Bellman operator ${{\mathcal{T}}_{B}}$. We note, however, that this optimality is achieved with great reduction in the sample complexity due in part to another important difference with standard RL, namely the search of our CBRL approach is across all states $x\in{{\mathbb{X}}}$ to find a single optimal parameter vector $p$ that derives a single optimal control-policy function ${\mathcal{G}}\in{\mathbb{F}}$ which coincides with the Bellman equation for each state. Iterations w.r.t. our CBRL operator ${{\mathbf{T}}}$ then consist of improving the estimates of the parameter vector $p$ while applying the optimal control-policy function ${\mathcal{G}}$ derived from the current estimate of $p$.
@@
Before addressing this iterative process within the context of policy gradient methods, we first relax some of the above conditions to consider families ${\mathbb{F}}$ that do not satisfy Assumption 1 and control-policy functions that map a parameter vector to a control policy whose domain spans across a large subset of states (as opposed to all states) in ${{\mathbb{X}}}$ and renders asymptotically optimal (as opposed to globally optimal) rewards. Supposing ${\mathbb{F}}$ satisfies Assumption 1, consider a sequence of less rich families ${\mathbb{F}}_{1}\subset\cdots\subset{\mathbb{F}}_{k}\subset{\mathbb{F}}$ of policy functions ${\mathcal{G}}^{(i)}\in{\mathbb{F}}_{i}$ derived from parameter vectors of the corresponding parameter sets ${{\mathbb{S}}}_{i}$, and define the operators ${{\mathbf{T}}}_{i}:\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})\rightarrow\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})$ as in (CBRL Approach) for any function $\tilde{q}_{i}(x,{\mathcal{F}}_{p}^{(i)})\in\tilde{{{\mathbb{Q}}}}({{\mathbb{X}}}\times\Pi_{i})$, $i\in[k]:=\{1,\ldots,k\}$. Then, from Theorem 1, each operator ${{\mathbf{T}}}_{i}$ under parameter set ${{\mathbb{S}}}_{i}$ is a contraction in the supremum norm and converges to the unique fixed point $\tilde{q}_{i}^{\prime}(x,{\mathcal{F}}_{p}^{(i)})$ that satisfies the corresponding version of (4), for all $x\in{{\mathbb{X}}}$ and $p\in{{\mathbb{S}}}_{i}$, $i\in[k]$. We then have the following formal result on the asymptotic optimality of our CBRL approach in such approximate settings.
Theorem 2.
@@
One specific instance of a sequence of the families of policy functions ${\mathbb{F}}_{1}\subset\cdots\subset{\mathbb{F}}_{k-1}\subset{\mathbb{F}}_{k}$ in Theorem 2 consists of piecewise-linear control policies of increasing richness (e.g., the class of canonical piecewise-linear functions (Lin and Unbehauen 1992)) w.r.t. finer and finer granularity of the control policy function space converging towards ${\mathbb{F}}$.
@@
Control-Policy-Parameter Gradient Ascent. We now return to our CBRL iterative process in the context of policy gradient methods, building on our foregoing results to establish theoretical results analogous to the standard policy gradient theorem that establishes, for policies parameterized by $\theta$, a connection between the gradient of the value function w.r.t. $\theta$ and the gradient of the policy action w.r.t. $\theta$. This theoretical result does not apply to our CBRL approach because there is no direct connection between the parameter vector $p\in{{\mathbb{S}}}$ and the control policy action. Rather, under CBRL and for any vector $p$, the control-policy function ${\mathcal{G}}(p)\in{\mathbb{F}}$ identifies a particular control policy ${\mathcal{F}}_{p}\in\Pi$, and then the policy applied in any state $x\in{{\mathbb{X}}}$ yields the action taken.